Why is a 7.0 earthquake ten times stronger than a 6.0 earthquake?
Found 2 solutions by stanbon, tsmith5651:
Answer by stanbon(75887)

Why is a 7.0 earthquake ten times stronger than a 6.0 earthquake?
The 1st earthquake has magnitude 10^7 I(o) where I(o) is the smallest
magnitude earthquake detectable by the Richter Scale.
The 2nd earthquake has magnitude 10^6 I(o) on the Richter Scale.
Note that one earthquake is ten time more powerful than the other.
Chers,
Stan H.

Answer by tsmith5651(7)

Lets talk about the Richter Scale. You know what the Richter Scale is, it's the way we measure the intensity of an earthquake. And if you're a local to Southern California, you know very well what a magnitude 3.3 or 6.4 earthquake means. The numbers on the Richter Scale are logarithmic. What that means is that as you increase by one integer, you are actually describing an increase in earthquake intensity by a factor of 10. In other words, a 6.0 earthquake is 10 times more intense than a 5.0 earthquake. You see, by increasing on the Richter Scale just by one, from 5.0 to 6.0, the 6.0 represents a ten-times increase in the force of the earthquake. A 7.0 earthquake is 10 times more intense than a 6.0 quake.
What if we compare a 7.0 quake and a 5.0 aftershock? How much weaker is the 5.0 aftershock? Well, from 7.0 to 6.0 means it is 10 times weaker. And when we go from 6.0 to 5.0, this is ten times weaker. So overall, if we go from 7.0 to 6.0, and then 6.0 to 5.0, we have represented a 10 times decrease multiplied by another 10 times decrease, or a (10x10)=100 times less intense aftershock.
Let's try another example. Is a quake registering 4.0 on the Richter Scale:
10 times stronger than a 2.0 quake?
100 times stronger than a 2.0 quake?
1000 times stronger than a 2.0 quake?
Let's simply figure it out. A 3.0 is ten times stronger than a 2.0 quake. And a 4.0 quake is 10 times stronger than the 3.0 quake. So when we move two digits up the Richter Scale, we're multipling 10 times stonger to another 10 times stronger. So
from 2.0 - 3.0 = 10 times stronger
from 3.0 - 4.0 = 10 times stronger
overall from 2.0 - 4.0 = (10 X 10) = 100 times stronger.
But having been 'shook, rattled and rolled' around by an earthquake and its aftershocks, you probably understand this idea quite well.